An air-dry soil with a moisture content of 25% by weight is to be used in a pot experiment. If each pot is to have 30kg oven-dried soil, how much of the air-dry soil should be weighed into the pot and how much of that moisture in kg?
1
Expert's answer
2021-10-15T12:32:20-0400

"m_w" =water mass

"m_t" = total mass

"(30+m_w) = m_t \\

m_w = 0.25m_t \\

30 + 0.25m_t = m_t \\

30 = 0.75m_t \\

m_t = 40 \;kg"

40 kg of the air-dry soil should be weighed into the pot.

"m_w = 0.25 \times 40 = 10 \;kg"

10 kg of that is moist